I am totally confused by the factoring results my text says is the answer: The equation is
3x^2 + x - 2 = 0
They say the factor of this is
(3x - 2)(x + 1)
Why isn't it -1, +2? Isn't the product of -2 suppose to equat the sum of 1 ?
Please show me why their answer is correct? I even tried using the ac Method with no results.

Answer by ewatrrr(24785)


Hi
3x^2 + x - 2 = 0 |After factoring, the factors must represent the original
factoring to a product (__ ± __)( ± __) Note: 2*-1 = -2 or -2*1 = -2
(3x ± __)(x ± __) |3x and x need to be in the 1st spots
(3x ± __)(x ± __) | 2 and 1 need to go into one or the other of 2nd spot blanks
(3x -2)(x+1)=0 |this arrangement works for the + and - using the 2 and 1
Noting: The SUM of the inner product(-2x) and the outer product(3x) = x
CHECKING our Answer
3x(x+1) - 2(x+1) = 3x^2 + 3x - 2x - 2 = 3x^2 + x - 2
(3x -2)(x+1)=0 | x = 2/3 or x = -1